# Running in circles: is practical application feasible for data fission and data thinning in post-clustering differential analysis?

### v1

@@ Abstract
The standard pipeline to analyse single-cell RNA sequencing (scRNA-seq) often involves two steps : clustering and Differential Expression Analysis (DEA) to annotate cell populations based on gene expression. However, using clustering results for data-driven hypothesis formulation compromises statistical properties, especially Type I error control. Data fission was introduced to split the information contained in each observation into two independent parts that can be used for clustering and testing. However, data fission was originally designed for non-mixture distributions, and adapting it for mixtures requires knowledge of the unknown clustering structure to estimate component-specific scale parameters. As components are typically unavailable in practice, scale parameter estimators often exhibit bias. We explicitly quantify how this bias affects subsequent post-clustering differential analysis Type I error rate despite employing data fission. In response, we propose a novel approach that involves modeling each observation as a realization of its distribution, with scale parameters estimated non-parametrically. Simulations study showcase the efficacy of our method when component are clearly separated. However, the level of separability required to reach good performance presents complexities in its application to real scRNA-seq data.
@@ 1 Introduction
Clustering encompasses all unsupervised statistical methods that group observations into homogeneous and separated clusters. Widely used in various application fields such as biology or genomics (Jaeger & Banks, 2023), clustering plays a significant role to uncover or summarise signals contained in different kind of multivariate data. In the context of single-cell RNA-seq (scRNA-seq) technologies which are high-throughput genomics techniques that measure gene expression at the single-cell level, providing insights into cellular heterogeneity and functional diversity within complex biological tissues, cluster analysis is an integral component of the traditional pipeline for data analysis (Amezquita et al., 2020). Combined with differential expression analysis, it allows to identify and annotate cellular subpopulations by simply grouping cells based on their gene expression. Differential expression analysis tests each individual gene for differential expression between two clusters, leading to the identification of marker genes that could potentially serve as cell-type marker genes (Pullin & McCarthy, 2024).
@@
However, such a two-step pipeline for post-clustering differential analysis requires using the same data twice: first to estimate the clusters, and then again to estimate the differences between them and perform significance testing – a procedure sometimes refered to as “double-dipping” (Kriegeskorte et al., 2009). In post-clustering differential analysis, it has been shown that the primary risk of double dipping is to compromise the control of the Type I error rate of otherwise well-calibrated testing procedure for traditional differential analysis, leading to false positives (Zhang et al., 2019). Fundamentally, uncertainties stemming from cluster analysis outcomes, particularly related with determining the optimal number of clusters, could create artificial differences between homogeneous groups of observations. Failure to account for the double use of the data during the analysis may lead differential analysis tools to erroneously identify those artificial differences (Hivert, Agniel, Thiébaut & Hejblum, 2024). Although challenges related to double dipping are increasingly studied, emphasizing their critical importance, they remain unresolved in the context of scRNA-seq data analysis (Lähnemann et al., 2020).
@@
Leveraging the selective inference framework (Fithian et al., 2014), which involves conditioning on the clustering event during the derivation of the test statistic and the associated p-value, various methodological efforts have been carried to address this double dipping issue in post-clustering inference (Gao et al., 2022). Zhang et al. (2019) introduced the Truncated-Normal test (TN-test), a four-step post-clustering differential analysis procedure that involves splitting the data into two parts, followed by cluster analysis on the first part of the data. Subsequently, a support vector machine (SVM) classifier is applied to the clustered data to learn the clustering structure and predict the cluster labels on the remaining data. Finally, a differential analysis is conducted between the predicted clusters using a truncated-normal test. This truncation, on either side of the hyperplane fitted by the SVM, allows to correct for the double dipping issue. However, information loss may occur as an inherent consequence of the data-splitting process. Moreover, the method involves multiple steps, introducing complexity into the overall analytical framework, and Song et al. (2023) highlighted the poor performances of the TN-test in their numerical simulations. More recently, Hivert, Agniel, Thiébaut & Hejblum (2024) and Chen & Gao (2023) have introduced selective tests tailored to detect mean differences between two clusters on individual variables. These methods explicitly condition on the clustering event to derive their p-value, relying on the set of all perturbed data sets that would yield the same partition when subjected to the same clustering algorithm. Hivert, Agniel, Thiébaut & Hejblum (2024) rely on a a Monte Carlo approach to suit any clustering algorithm, which comes at the expense of extensive computational times. On the other hand, Chen & Gao (2023) explicitly describe this set specifically for K-means and hierarchical clustering applied to the squared distance matrix. Bachoc et al. (2023) introduced a selective tests designed for convex clustering. All these proposed selective methods designed for post-clustering inference either require specific clustering algorithms or introduce new specific test statistics. This consequently increases their complexity and makes their application less straightforward in the context of scRNA-seq data analysis.
@@
Leiner et al. (2023) have drawn inspiration from data splitting (which effectively addresses overfitting issues in machine learning) to break free from the selective inference framework. They propose a method called “data fission”, wherein the information within each individual observation is split into two independent parts. The first part could be used for cluster analysis, and labeling observations on the second part simply by matching them to their first counterpart. Differential analysis could then be conducted on the remaining information, i.e. the second part. However, in the context of post-clustering differential analysis, it is imperative that the two parts be independent, as each analysis (cluster analysis and differential analysis) must be performed independently to effectively prevent double dipping and the associated inflation of Type I error. Data fission requires strong parametric distributional assumptions, with only the Gaussian and Poisson (Neufeld et al., 2024) distributions ensuring independence between the two fissioned parts. Expanding upon this concept, Neufeld, Dharamshi, Gao & Witten (2023) have generalized the fission process by introducing “data thinning”. Building on the same foundational idea, they not only succeed in developing a process capable of decomposing data into more than two parts, but also broaden the spectrum of distributions where independence between each part is provided. This includes the negative binomial distribution, widely used when modeling RNA-seq data.
@@
Data fission or data thinning, although initially appealing, present significant limitations. First, these methods lack results and justification when applied to mixture distributions, which are commonly used to model data with a clustered structure (MacQueen et al., 1967). Consequently, the absence of such results inherently assumes a global null hypothesis of no clusters in the data. Additionally, these methods assume prior knowledge of parameters (e.g., variance for the Gaussian distribution or overdispersion for the negative binomial distribution). Although robust estimators for these parameters could theoretically ensure the validity of the method, this further adds complexity for clustered data where each cluster has a different parameter value. And in the absence of knowledge about the data structure, specifically the clusters, the only justifiable estimator is the full-sample one (i.e. computed using all the observations regardless of their mixture component), even though it fails to correctly estimate the intra-component parameter value.
@@
Data fission and data thinning have emerged as promising alternatives to selective inference for post-clustering inference, owing to their compatibility with any clustering methods and differential analysis tests. In this article, we propose a strategy to address their inherent limitations by modeling each observation as a realization from distinct random variables with their own parameter value. This approach facilitates the application of fission or thinning processes within the context of mixture distributions although the accurate estimation of scale parameters still remains challenging. We establish a link between the bias in the estimation of the variance parameter in the Gaussian distribution and the expected Type I error rate of the two-sample $t$-test (Welch, 1947), underscoring the critical importance of a robust estimator and the practical complexities associated with implementing these methods. To tackle this issue, we introduce a non-parametric method for estimating the local individual variance to be used in the Gaussian setting.
@@ 2.1 Data fission and data thinning
Let $\boldsymbol{X}$ be a random variable with a known distribution. Both data fission (Leiner et al., 2023) and its generalization, data thinning (Neufeld, Dharamshi, Gao & Witten, 2023), aim to decompose the random variable $\boldsymbol{X}$ into two (or more in the case of data thinning) new random variables $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$. These new variables are designed to i) retain information from the original variable $\boldsymbol{X}$, and ii) be independent. Of note, data fission can generate pairs of $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ that are not independent, but her we will focus on the independent cases only. The balance between the amount of information from $\boldsymbol{X}$ kept into either $\boldsymbol{X}^{(1)}$ or $\boldsymbol{X}^{(2)}$ is tuned by a hyperparameter $\tau$. Such a decomposition can be performed for various probability distributions of $\boldsymbol{X}$. For data fission, Leiner et al. (2023) identified only two distributions, the Gaussian and the Poisson, that satisfy the independence requirement. In contrast, Neufeld, Dharamshi, Gao & Witten (2023) established a comprehensive decomposition that ensures independence for all convolution-closed distributions. This encompasses Gaussian, Poisson (Neufeld et al., 2024) and Negative Binomial (Neufeld, Popp, Gao, Battle & Witten, 2023) distributions.
@@
$\mathcal{N}(\boldsymbol{\mu}_{p},\boldsymbol{\Sigma}_{p\times p})$ $\tau\in]0,+\infty)$
@@
$\boldsymbol{X}^{(1)}\,|\,\boldsymbol{X}=\boldsymbol{x}\sim\mathcal{N}(\tau\boldsymbol{x},\tau(1-\tau)\boldsymbol{\Sigma}_{p\times p})$
@@
$\mathcal{NB}(\mu,\theta)$ $\tau\in[0,1]$ No fission
@@
Decompositions for the three distributions into two independent random variables $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$, are detailed in Table 1. For proofs of independence, please refer to the Appendix A for the Gaussian case, and to Neufeld et al. (2024) and Neufeld, Popp, Gao, Battle & Witten (2023) for the Poisson and negative binomial distribution, respectively. Both Gaussian and negative binomial data fissions/thinnings require knowledge of scale parameters (namely $\boldsymbol{\Sigma}$ or $\theta$) for practical feasibility, as highlighted in Table 1. Theoretical guarantees, and especially the independence between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$, are based on using the true values of these parameters. Yet, in real-life applications these are unknown and need to be estimated, most likely from the data as described in Neufeld, Popp, Gao, Battle & Witten (2023).
2.2 Limits in practical application of data fission an data thinning
@@ 2.2.1 Mixture distributions
This can thus only be verified for $K=1$, so under the global null of no cluster. Yun & Foygel Barber (2023) highlighted similar challenges for existing post-clustering selective tests.
@@ 2.2.2 Scale parameter prior knowledge and estimation
Figure 2: Toy example illustrating the impact of variance estimation on data fission p-values Panel A: A two-dimensional Gaussian distribution incorrectly clustered into $2$ clusters. Panel B: QQ-plot of the $t$-test p-values for the comparison between the two estimated clusters across $1,\!000$ simulations when data fission is performed with $3$ different variance estimators. Panel C: Extension of the problem to a two-component, two-dimensional Gaussian mixture incorrectly clustered into $3$ clusters. $C_{1}$ and $C_{2}$ originate from the same component, which is erroneously split into two. Panel D: $t$-test p-values for the comparison between $C_{1}$ and $C_{2}$ over $1,\!000$ data simulations using the same $3$ variance estimators for the data fission.
@@
Figure 2 provides a comprehensive overview of the challenges associated with variance estimation in data fission. Panel A presents an illustrative example involving $300$ realizations of a multivariate Gaussian distribution ($K=1$) with a mean vector $\boldsymbol{\mu}=\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right)$ and a covariance matrix $\boldsymbol{\Sigma}=\left(\begin{smallmatrix}4&0\\ 0&1\end{smallmatrix}\right)$. The k-means algorithm was applied to this dataset, resulting in the estimation of two clusters, $C_{1}$ and $C_{2}$. Data fission performance was evaluated when using different estimations of $\boldsymbol{\Sigma}$. First, we considered the true intra-components covariance matrix $\boldsymbol{\Sigma}_{k}$ (see section 2.3.1 for a proposed approach to data fission assuming several mixture component). We also considered the overall sample covariance matrix $\widehat{\boldsymbol{\Sigma}}=\frac{1}{n-1}\sum_{i=1}^{n}\left(\boldsymbol{x}_{i}-\overline{\boldsymbol{X}}\right)\left(\boldsymbol{x}_{i}-\overline{\boldsymbol{X}}\right)^{t}$ where $\overline{\boldsymbol{X}}$ is the sample mean vector. We finally use the $k-$means results to compute an intra-cluster covariance matrix defined as: $\widehat{\boldsymbol{\Sigma}}_{\hat{k}}=\frac{1}{|C_{k}|-1}\sum_{i=1\in C_{k}}\left(\boldsymbol{x}_{i}-\overline{\boldsymbol{X}}_{C_{k}}\right)\left(\boldsymbol{x}_{i}-\overline{\boldsymbol{X}}_{C_{k}}\right)^{t}$ where $\overline{\boldsymbol{X}}_{C_{k}}$ is the sample mean vector of the cluster $C_{k}$. Since both clusters originate from the same component, there is no inherent differences between them, implying that the $t$-test p-values should exhibit a uniform distribution. Panel B presents a QQ-plot illustrating the resulting p-values against the Uniform distribution for the test on $X_{1}$ in $1,\!000$ replications of the experiment. In this scenario where the mixture has only one component ($K=1$), the sample covariance matrix $\widehat{\boldsymbol{\Sigma}}$ provides an unbiased estimate of the true $\boldsymbol{\Sigma}$, resulting in uniformly distributed p-values. However, when considering the intra-cluster covariance matrices $\widehat{\boldsymbol{\Sigma}}_{\hat{k}}$ (for $k=1,2$) derived from the k-means results, the p-values no longer exhibit a uniform distribution. In this case, the estimated matrices $\widehat{\boldsymbol{\Sigma}}_{\hat{k}}$ for each cluster drastically underestimate the true covariance matrix $\boldsymbol{\Sigma}$, compromising the independence between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$. This deviation from independence indicates that it becomes straightforward to replicate the clusters identified in $\boldsymbol{X}^{(1)}$ onto $\boldsymbol{X}^{(2)}$.
@@
Panel C introduces a scenario with two true clusters generated using a mixture of two Gaussian distributions ($K=2$). The K-means identifies $3$ clusters (for illustration purposes), incorrectly splitting one mixture component into clusters $C_{1}$ and $C_{2}$. We compared data fission performance was again compared when the decomposition is performed using the same $3$ covariance estimators as in the previous scenario. Panel D presents the resulting p-values for the $t$-test on $C_{1}$ and $C_{2}$ on $X_{1}$ over $1,\!000$ replications of the analysis. Efficient Type I error control is achievable only by considering the true intra-components covariance matrix. Both global and estimated intra-cluster covariances are biased estimators, leading to correlations between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$, compromising control of the Type I error rate. This underscores the intricate challenges associated with covariance estimation for data fission in practical scenarios, mainly due to the misspecification of the generative model for data fission arising when genuine clusters exist within the dataset.

### v2

# Running in circles: practical limitations for real-life application of data fission and data thinning in post-clustering differential analysis

@@ Abstract
Post-clustering inference in scRNA-seq analysis presents significant challenges in controlling Type I error during Differential Expression Analysis. Data fission, a promising approach, aims to split the data into two new independent parts, but relies on strong parametric assumptions of non-mixture distributions, which are violated in clustered data. We show that applying data fission to these mixtures requires knowledge of the clustering structure to accurately estimate component-specific scale parameters. These estimates are critical for ensuring decomposition and independence. We theoretically quantify the direct impact of the bias in estimating this scales parameters on the inflation of the Type I error rate, caused by a deviation from the independence. Since component structures are unknown in practice, we propose a heteroscedastic model with non-parametric estimators for individual scale parameters. This model uses proximity between observations to capture the effect of the underlying mixture on data dispersion. While this approach works well when clusters are well-separated, it introduces bias when separation is weak, highlighting the difficulty of applying data fission in real-world scenarios with unknown degrees of separation.
@@ 1 Introduction
Clustering encompasses all unsupervised statistical methods that group observations into homogeneous and separated clusters. Widely used in various application fields such as biology or genomics (Jaeger and Banks,, 2023), clustering plays a significant role to uncover or summarize signals contained in different kind of multivariate data. In the context of single-cell RNA-seq (scRNA-seq) technologies which are high-throughput genomics techniques that measure gene expression at the single-cell level, providing insights into cellular heterogeneity and functional diversity within complex biological tissues, cluster analysis is the first step of the traditional pipeline for data analysis (Amezquita et al.,, 2020). The clustering groups cells based on their gene expression/abundance. Afterwards, the differential expression analysis, comparing gene abundance between groups, allows to identify and annotate cellular sub-populations. This leads to the identification of marker genes that could potentially serve as cell-type marker genes (Pullin and McCarthy,, 2024).
@@
However, such a two-step pipeline for post-clustering differential analysis requires using the same data twice: first to estimate the clusters, and then again to estimate the differences between them and perform significance testing – a procedure sometimes referred to as “double-dipping” (Kriegeskorte et al.,, 2009). In post-clustering differential analysis, it has been shown that the primary risk of double dipping is to compromise the control of the Type I error rate of otherwise well-calibrated testing procedure for traditional differential analysis, leading to false positives (Zhang et al.,, 2019). Fundamentally, uncertainties stemming from cluster analysis outcomes, particularly related with determining the optimal number of clusters, could create artificial differences between homogeneous groups of observations. Failure to account for the double use of the data during the analysis may lead differential analysis tools to erroneously identify those artificial differences (Hivert et al., 2024a, ). Although challenges related to double dipping are increasingly studied, they remain unresolved in the context of scRNA-seq data analysis (Lähnemann et al.,, 2020). We described below the proposed approaches and the remaining challenges.
@@
Leveraging the selective inference framework (Fithian et al.,, 2014), which involves conditioning on the clustering event during the derivation of the test statistic and the associated p-value, various methodological efforts have been proposed to address this double dipping issue in post-clustering inference (Gao et al.,, 2022). Zhang et al., (2019) introduced the Truncated-Normal test (TN-test), a four-step post-clustering differential analysis procedure that involves splitting the data into two parts, followed by cluster analysis on the first part of the data. Subsequently, a support vector machine (SVM) classifier is applied to the clustered data to learn the clustering structure and predict the cluster labels on the remaining data. Finally, a differential analysis is conducted between the predicted clusters using a truncated-normal test. This truncation, on either side of the hyperplane fitted by the SVM, allows to correct for the double dipping issue. However, information loss may occur as an inherent consequence of the data-splitting process. Moreover, the method involves multiple steps, introducing complexity into the overall analytical framework, and Song et al., (2023) highlighted the poor performances of the TN-test in their numerical simulations. More recently, Hivert et al., 2024a and Chen and Gao, (2023) have introduced univariate selective tests tailored to detect mean differences between two (multivariate) clusters. These methods explicitly condition on the clustering event to derive their p-value, relying on the set of all perturbed data sets that would yield the same partition when subjected to the same clustering algorithm. Hivert et al., 2024a rely on a Monte Carlo approach to suit any clustering algorithm, which comes at the expense of extensive computational times. On the other hand, Chen and Gao, (2023) explicitly describe this set specifically for $k-$means and hierarchical clustering applied to the squared distance matrix. Bachoc et al., (2023) also introduced a selective test designed for convex clustering. All these proposed selective methods designed for post-clustering inference either require specific clustering algorithms or introduce new specific test statistics. This consequently increases their complexity and makes their application less straightforward in the context of scRNA-seq data analysis.
@@
Leiner et al., (2023) have drawn inspiration from data splitting (which effectively addresses overfitting issues in machine learning) to break free from the selective inference framework. They propose a method called “data fission”, wherein the information within each individual observation is split into two independent parts. The first part could be used for cluster analysis, and labeling observations on the second part simply by matching them to their first counterpart. Differential analysis could then be conducted on the remaining information, i.e. the second part. However, in the context of post-clustering differential analysis, it is imperative that the two parts be independent, as each analysis (cluster analysis and differential analysis) must be performed independently to effectively prevent double dipping and the associated inflation of Type I error. Data fission requires strong parametric distributional assumptions, with only the Gaussian and Poisson (Neufeld et al.,, 2024) distributions ensuring independence between the two fissioned parts. Expanding upon this concept, Neufeld et al., 2023a have generalized the fission process by introducing “data thinning”. Building on the same foundational idea, they not only succeed in developing a process capable of decomposing data into more than two parts, but also broaden the spectrum of distributions where independence between each part is provided. This includes the negative binomial distribution, widely used when modeling RNA-seq data.
@@
Data fission and data thinning have emerged as promising alternatives to selective inference for post-clustering inference due to their compatibility with various clustering methods and differential analysis tests. However, they present some inherent limitations that make them difficult to apply for post-clustering inference. First, these methods lack results and justification when applied to mixture distributions, which are commonly used to model data with a clustered structure (Macqueen,, 1967). Consequently, the absence of such results inherently assumes a global null hypothesis of no clusters in the data when applying data fission or data thinning. Additionally, these methods assume prior knowledge of parameters (e.g., variance for the Gaussian distribution or overdispersion for the negative binomial distribution). Although robust estimators for these parameters could theoretically ensure the validity of the method, this further adds complexity for clustered data where each cluster has a different parameter value. In the absence of knowledge about the data structure, specifically the clusters, the only justifiable estimator is the full-sample one (i.e. computed using all the observations regardless of their mixture component) that fails to correctly estimate the intra-component parameter value.
@@
In this article, we demonstrate that these approaches are not practical for real-world applications. Performing cluster analysis inherently assumes that the data originate from mixture models, contradicting the parametric assumptions of data fission and data thinning, which cannot decompose such distributions. Even if it is possible to move beyond the framework of mixture models by modeling each observation as a realization from distinct random variables with their own parameter values, accurately estimating scale parameters remains challenging without knowledge of the mixture. Specifically, we establish a link between the bias in estimating the variance parameter in the Gaussian distribution and the expected Type I error rate of the two-sample $t$-test (Welch,, 1947), underscoring the critical importance of a robust estimator. We employ a non-parametric method for estimating local variance in the Gaussian setting to try to use proximity between observations as a proxy of the unknown underlying mixture. However the poor performances of this approach demonstrates that accurate variance estimation, and hence control of the Type I error rate, remains challenging without knowing the mixture, i.e. the clustering, prior to decomposition.
@@ 2.1 Data fission and data thinning
Data fission and data thinning are methods designed to decompose a random variable into two new independent random variables that can ensure the post-clustering inference validity.
Let $\boldsymbol{X}$ be a random variable with a known distribution. Both data fission (Leiner et al.,, 2023) and its generalization, data thinning (Neufeld et al., 2023a, ), aim to decompose the random variable $\boldsymbol{X}$ into two (or more in data thinning) new random variables $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$. These new variables are designed to i) retain information from the original variable $\boldsymbol{X}$, and ii) be independent. Of note, data fission can generate pairs of $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ that are not independent, but her we will focus on the independent cases only. The balance between the amount of information from $\boldsymbol{X}$ kept in either $\boldsymbol{X}^{(1)}$ or $\boldsymbol{X}^{(2)}$ is tuned by a hyperparameter $\tau$. Such a decomposition can be performed for various probability distributions of $\boldsymbol{X}$. For data fission, Leiner et al., (2023) identified only two distributions, the Gaussian and the Poisson, that satisfy the independence requirement. In contrast, Neufeld et al., 2023a established a comprehensive decomposition that ensures independence for all convolution-closed distributions. This encompasses Gaussian, Poisson (Neufeld et al.,, 2024) and Negative Binomial (Neufeld et al., 2023b, ) distributions.
@@
$\mathcal{N}(\boldsymbol{\mu}_{p},\boldsymbol{\Sigma}_{p\times p})$
$\tau\in]0,+\infty)$
$\tau_{2}\in]0,1[$
@@
$\boldsymbol{X}^{(1)}\,|\,\boldsymbol{X}=\boldsymbol{x}\sim\mathcal{N}(\tau_{2}\boldsymbol{x},\tau_{2}(1-\tau_{2})\boldsymbol{\Sigma}_{p\times p})$
@@
$\operatorname{NegBin}(\mu,\theta)$ $\tau\in[0,1]$ No fission
@@
Decompositions for the three distributions into two independent random variables $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$, are detailed in Table 1. For proofs of independence, please refer to the Section $1$ of the Supplementary Materials for the Gaussian case, and to Neufeld et al., (2024) and Neufeld et al., 2023b for the Poisson and negative binomial distribution, respectively. Both Gaussian and negative binomial data fissions/thinnings require knowledge of scale parameters (namely $\boldsymbol{\Sigma}$ or $\theta$) for practical feasibility, as highlighted in Table 1. Theoretical guarantees, and especially the independence between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$, are based on using the true values of these parameters. Yet, in real-life applications these are unknown and need to be estimated, most likely from the data as described in Neufeld et al., 2023b .
2.2 Limits in practical application of data fission and data thinning
Data fission and data thinning methods face a circular challenge when applied for post-clustering inference to mixture distributions used to describe clustered data. They require accurate estimates of intra-component parameters (like the variance), which depend on knowing the true cluster – but estimating the clusters is the whole point of the analysis in the first place, and true clusters are never known a priori.
@@ 2.2.1 Mixture distributions
This corresponds to the global null of no cluster. Yun and Foygel Barber, (2023) highlighted similar challenges for existing post-clustering selective tests.
@@ 2.2.2 Scale parameter prior knowledge and estimation
Figure 2: Toy example illustrating the impact of variance estimation on data fission p-values Panel A: A two-dimensional Gaussian distribution incorrectly clustered into $2$ clusters. Panel B: QQ-plot of the $t$-test p-values for the comparison between the two estimated clusters across $1,\!000$ simulations when data fission is performed with $3$ different variance estimators. Panel C: Extension of the problem to a two-component, two-dimensional Gaussian mixture incorrectly clustered into $3$ clusters. $C_{1}$ and $C_{2}$ originate from the same component, which is erroneously split into two. Panel D: $t$-test p-values for the comparison between $C_{1}$ and $C_{2}$ over $1,\!000$ data simulations using the same three variance estimators for the data fission.
@@
Figure 2 provides a comprehensive overview of the challenges associated with variance estimation in data fission. Panel A presents an illustrative example involving $300$ realizations of a multivariate Gaussian distribution ($K=1$) with a mean vector $\boldsymbol{\mu}=\left(\begin{smallmatrix}0\\ 0\end{smallmatrix}\right)$ and a covariance matrix $\boldsymbol{\Sigma}=\left(\begin{smallmatrix}4&0\\ 0&1\end{smallmatrix}\right)$. The $k-$means algorithm was applied to this dataset, resulting in the estimation of two clusters, $C_{1}$ and $C_{2}$. Data fission performance was evaluated when using different estimations of $\boldsymbol{\Sigma}$. First, we considered the true intra-components covariance matrix $\boldsymbol{\Sigma}_{k}$ (see section 2.3.1 for a proposed approach to data fission assuming several mixture component). We also considered the overall sample covariance matrix $\widehat{\boldsymbol{\Sigma}}=\frac{1}{n-1}\sum_{i=1}^{n}\left(\boldsymbol{x}_{i}-\overline{\boldsymbol{X}}\right)\left(\boldsymbol{x}_{i}-\overline{\boldsymbol{X}}\right)^{t}$ where $\overline{\boldsymbol{X}}$ is the sample mean vector. We finally use the $k-$means results to compute an intra-cluster covariance matrix defined as: $\widehat{\boldsymbol{\Sigma}}_{\hat{k}}=\frac{1}{|C_{k}|-1}\sum_{i=1\in C_{k}}\left(\boldsymbol{x}_{i}-\overline{\boldsymbol{X}}_{C_{k}}\right)\left(\boldsymbol{x}_{i}-\overline{\boldsymbol{X}}_{C_{k}}\right)^{t}$ where $\overline{\boldsymbol{X}}_{C_{k}}$ is the sample mean vector of the cluster $C_{k}$. Since both clusters originate from the same component, there is no inherent differences between them, implying that the $t$-test p-values should exhibit a uniform distribution. Panel B presents a QQ-plot illustrating the resulting p-values against the Uniform distribution for the test on $X_{1}$ in $1,\!000$ replications of the experiment. In this scenario where the mixture has only one component ($K=1$), the sample covariance matrix $\widehat{\boldsymbol{\Sigma}}$ provides an unbiased estimate of the true $\boldsymbol{\Sigma}$, resulting in uniformly distributed p-values. However, when considering the intra-cluster covariance matrices $\widehat{\boldsymbol{\Sigma}}_{\hat{k}}$ (for $k=1,2$) derived from the $k-$means results, the p-values no longer exhibit a uniform distribution. In this case, the estimated matrices $\widehat{\boldsymbol{\Sigma}}_{\hat{k}}$ for each cluster drastically underestimate the true covariance matrix $\boldsymbol{\Sigma}$, compromising the independence between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$. This deviation from independence leads to spuriously concluding the estimated clusters are truly different from one another.
@@
Panel C introduces a scenario with two true clusters generated using a mixture of two Gaussian distributions ($K=2$). The $k-$means identifies $3$ clusters (for illustration purposes), incorrectly splitting one mixture component into clusters $C_{1}$ and $C_{2}$. Data fission performance was again compared when the decomposition is performed using the same $3$ covariance estimators as in the previous scenario. Panel D presents the resulting p-values for the $t$-test on $C_{1}$ and $C_{2}$ on $X_{1}$ over $1,\!000$ replications of the analysis. Efficient Type I error control is achievable only by considering the true intra-components covariance matrix. Both global and estimated intra-cluster covariances are biased estimators, leading to correlations between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$, compromising control of the Type I error rate. This underscores the intricate challenges associated with covariance estimation for data fission in practical scenarios, mainly due to the misspecification of the generative model for data fission arising when genuine clusters exist within the dataset.